OpenStudy (partyshotz2409):

If x^2 = 10, what is the value of x? ±square root of 10 ±square root of 20 ±5 ±20

OpenStudy (supersmart1001):

www.montereyinstitute.org/courses/DevelopmentalMath/COURSE_TEXT2_RESOURCE/U16_L5_T1_text_final.html im not sure if this will help but its the best I can do

jimthompson5910 (jim_thompson5910):

Similar Example: \[\Large x^2 = 70\] \[\Large \sqrt{x^2} = \sqrt{70}\] \[\Large |x| = \sqrt{70}\] \[\Large x = \pm\sqrt{70}\] Notes: For step 2, apply the square root to both sides. This will undo the squaring done to the 'x' For step 3, use the rule \(\Large \sqrt{x^2} = |x|\) where x is any real number For step 4, use the rule that if |x| = k, then x = -k or x = k where k is some positive number.

OpenStudy (partyshotz2409):

thnx